If 60 cubic meter of water are pumped per hour to a height of 80 meters calculate the power required in kilowatts
1
Expert's answer
2019-12-16T11:12:33-0500

First of all let's calculate the mass of water in 60 cubic meter

density of water = 1000kg/m3

mass of water = 60*1000 = 60000 kg

Energy= mgH where m being the mass, and H being the height

=60000*9.8*80

=47040000 joules

power =energy/time

=47040000/(60*60)

=13066.67 watt

so power in kilowatt = 13066.67/1000

=13.06667 kilowatt

=13.06 kilowatt